OpenStudy (anonymous):

integrate cosx.cos3x

Parth (parthkohli):

\[\cos(x) \cos(3x) = \frac{1}{2}(\cos(2x) + \cos (4x))\]

Parth (parthkohli):

So\[\int \dfrac{1}{2}(\cos(2x)+ \cos(4x))dx = \frac{1}{2}\int [\cos(2x) + \cos(4x)]dx \]

Parth (parthkohli):

The rest should be easy to handle.

OpenStudy (anonymous):

i got it .thanx